(-2)^2 -(3)^2/(1-4)^2x2-6=
Answer by stanbon(48535)

(-2)^2 -[(3)^2/(1-4)^2] x 2 - 6
= 4 - 9 / (9)x2 - 6
= 4 - (9/18) - 6
= 4 -(1/2) - 6
= -2 1/2
================
Cheers,
Stan H.